# Learning a Stable Dynamic System with a Lyapunov Energy Function for Demonstratives Using Neural Networks

### v4

@@ I INTRODUCTION
The popularity of robotics is growing with advancements in technology, and robots are expected to possess greater intelligence to perform complex tasks comparable to those performed by humans. However, programming them, especially for non-experts, is challenging. In such situations, Learning from Demonstration (LfD) is a user-friendly approach for robots to acquire skills by observing demonstrations, without relying on an explicit script or defined reward functions [1].
@@
While CLF-DM has the advantage of being able to learn from a broader set of demonstrations, it should be noted that the algorithm requires solving optimization problems in each step to maintain stability. This complexity and sensitivity to parameters are important considerations.
@@
To address the challenges commonly encountered by neural network-based algorithms, in this paper, a novel neural network-based approach is proposed that eliminates the need for invertible transformations. The key contributions of this paper are summarized as follows:
  • A novel neural network architecture is proposed that is capable of directly learning a Lyapunov candidate function from demonstration data.
@@
The paper is organized as follows: Section II provides a comprehensive discussion of the problem formulations. In section III, the intricate details of the neural networks are presented. In section IV, the evaluation results of the proposed algorithm. This evaluation includes simulations conducted on various handwriting trajectories sourced from the Lasa datasets [12], as well as experiments carried out on the Franka Emika robot. In section V, the conclusion is made.
@@ II Problem Formulation
where $f:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ is a continuous and continuously differentiable function, which is described to have a single equilibrium state, denoted as $\boldsymbol{x}^{*}$. This equilibrium state represents the target state of the task and can be viewed as the attractor of the DS. To simplify the analysis and without sacrificing generality, the targets of the motions are assumed to be positioned at the origin of the Cartesian coordinate system by $\tilde{\boldsymbol{x}}=\boldsymbol{x}-\boldsymbol{x}^{*}$. Equation (1), which can also describe the manipulation skills, yields a solution $\Phi(\boldsymbol{x}_{0},t)$ that represents a valid trajectory generated by the Autonomous DS when provided with an initial state $\boldsymbol{x}_{0}$. The ideal motions can then be obtained by adding this trajectory to the target state $\boldsymbol{x}=\tilde{\boldsymbol{x}}+\boldsymbol{x}^{*}$. Consequently, by altering the initial conditions $\boldsymbol{x}_{0}$, different trajectories leading to the target state $\boldsymbol{x}^{*}$ can be generated. This flexibility allows for various trajectories to be produced, each tailored to different initial conditions.
The demonstration data is typically recorded in the following format: $\begin{Bmatrix}\boldsymbol{x}_{k,n},\dot{\boldsymbol{x}}_{k,n}\ |\ k=1,2,...,K_{n};\ n=1,2,...,N_{d}\end{Bmatrix},$ where $n$ represents the index of the demonstrations and $k$ is the $k$-th sampling time. $N_{d}$ denotes the total number of demonstrations and $K_{n}\ (n=1,2,...,N_{d})$ represents the whole sampling number of the $n$-th demonstrated motion. The state variable $\boldsymbol{x}$ signifies the condition or configuration of either a human or a robot. Typically, in the context of robotic systems, it corresponds to the spatial coordinates of the robot’s end-effector. Alternatively, in the case of human involvement, it characterizes the spatial position of the human hand within Cartesian space. $\dot{\boldsymbol{x}}$ is the first-order time derivative of $\boldsymbol{x}$. Moreover, it’s important to note that all the demonstrations for a given task share the same target state. It is assumed that these demonstrations conform to an Autonomous DS as described in Equation (1). This alignment allows for the modeling of the system in a parametric fashion, which can be represented as:
@@
where $\hat{\dot{\boldsymbol{x}}}$ denotes the estimate of $\dot{\boldsymbol{x}}$, $\hat{f}$ is the model designed manually to approximate the DS in (1), and $\boldsymbol{\theta}$ denotes the parameter in the model which can be learned. The optimal parameter configuration, denoted as $\boldsymbol{\theta^{*}}$, is obtained through the minimization of the following objective function:
@@
where $\propto$ denotes the proportionality relation, $\left\|\cdot\right\|^{2}$ is the $l_{2}$-norm and $\hat{\dot{\boldsymbol{x}}}_{k,n}$ represents the estimate of $\dot{\boldsymbol{x}}_{k,n}$. Therefore, the objective function compels the learned model to generate precise reproductions that closely match the demonstrated behavior. Nonetheless, it’s crucial to emphasize that the stability of the model cannot be guaranteed solely by the objective function. An Autonomous Dynamical System (DS) is deemed to be locally asymptotically stable at the point $\boldsymbol{x}^{*}$ if there exists a continuous and continuously differentiable Lyapunov candidate function $V(\boldsymbol{x}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ that satisfies the following conditions:
@@
In this paper, a data-driven approach to learn the Lyapunov candidate function, utilizing neural networks is proposed. This method offers an alternative that bypasses the need for designing Lyapunov candidate function, facilitating the achievement of stability in the DS learning process, which is introduced in the following section.
@@ III The Proposed Approach
Figure 2: The simulation employing the proposed algorithm is illustrated. The images with a dark background showcase the learned vector fields. Within these vector field visuals, the white dotted lines represent the original demonstration data, while the red solid lines correspond to the reproductions, all originating from the same initial points. The target points are denoted as ”$x$” in these visuals. Furthermore, the transformed reproduction trajectories (consistent with $\boldsymbol{y}$ as shown in Fig. 1) are presented with white background. Within this context, the solid points indicate the starting positions for these trajectories.
@@
$\dot{\boldsymbol{y}}=\begin{cases}n(\boldsymbol{y})\ \ \ \ \ \ if\ \boldsymbol{y}^{\mathrm{T}}n(\boldsymbol{y})<0,\\ -\beta\boldsymbol{y}\ \ \ \ \ \ \ \ \ otherwise,\end{cases}$ (13)
where $n(\cdot)$ is a neural network and $\beta$ is a manually designed positive constant. Taking inspiration from the research conducted by [13] and simultaneously learning both $\dot{\boldsymbol{y}}$ and $\dot{\boldsymbol{x}}$, the equation (12) can be reformulated as:
$\dot{\boldsymbol{y}}=n(\boldsymbol{y})-(n(\boldsymbol{y})+\beta\boldsymbol{y})\frac{\mathrm{Relu}(\boldsymbol{y}^{\mathrm{T}}n(\boldsymbol{y}))}{\boldsymbol{y}^{\mathrm{T}}n(\boldsymbol{y})},$ (14)
@@
To address the third problem, the neural network structure can be configured in a way that ensures the output becomes a zero vector only when the input is a zero vector. One simple way to enhance the performance of the neural network is to utilize a suitable activation function, such as the $\mathrm{Softplus}$ or $\mathrm{Sigmoid}$ function, which guarantee that the output is always greater than zero. Alternatively, the $\mathrm{Tanh}$ function can be employed to ensure that the output is always lesser than one. Furthermore, the output of the neural network should be multiplied with the input vector $\boldsymbol{x}$ and then added to it. In our paper, the expression for $\boldsymbol{y}$ corresponding with the illustration in Fig. 1 that is defined as follows:
@@
Interestingly, when there is a one-to-one correspondence between the target points in two spaces, the input is $\infty$ then the output is also $\infty$, and the two spaces have the same dimensionality, the transformation can be referred to as a bijective transformation. By contradiction, the two-dimensional space is first taken into consideration, and it is assumed that the transformation from $\boldsymbol{x}$ to $\boldsymbol{y}$ is not bijective. In this case, it can be deduced that two vectors $\boldsymbol{x}$ are transformed to the same vector $\boldsymbol{y}$, such as $g(\boldsymbol{x_{1}})=g(\boldsymbol{x_{2}})$. Since equation 7 sets the Lyapunov energy function, it follows that the energy of both $x_{1}$ and $x_{2}$ is the same. The points in the transformed 2-dimensional space, where all have the same energy, can be combined to form a closed and continuous circle. Similarly, in the original space, the points with the same energy must be combined to form a continuous curve $l$ that connects $x_{1}$ and $x_{2}$. Since the energy of the target point is zero and the energy of a point at infinity is infinite, a curve must exist that connects the point with the target point but does not intersect with curve $l$. It is not possible for the energy in the existing curve to be continuous. Thus the two-dimensional transformation must be a bijective transformation. When the transformation space is more than two-dimensional, it is possible for the projection space to also be two-dimensional, and in this case, the contradiction remains. Therefore, the transformation can be referred to as a bijective transformation.
@@ IV-A Simulation Results
To assess the effectiveness of the proposed algorithm, the quantitative evaluation of accuracy was performed through the application of two error metrics: the Swept Error Area (SEA) [5] and the Velocity Root Mean Square Error ($V_{rmse}$) [14]. The SEA scores serve as an indicator of the algorithm’s capability to faithfully replicate the shapes of the demonstrated motions. On the other hand, the $V_{rmse}$ metric gauges the algorithm’s proficiency in preserving the velocities of the demonstrations. A lower SEA score signifies superior accuracy in reproducing the trajectories with the same initial start points, while a lower $V_{rmse}$ value indicates that the reproductions closely match the smoothness of the original demonstrations. These metrics collectively assess the performance of the reproduced trajectories by quantifying their similarity to the demonstrated trajectories.
@@
For the implementation of the proposed algorithm, the ”AdamW” optimization method is opted, adhering to the default hyper-parameters. Specifically, the initial learning rate was set to $1\times 10^{-5}$, with a decay rate of $0.99$. Subsequently, the learned DS was utilized to generate trajectories originating from the same starting points with the same steps as demonstrations. In this paper, all demonstration data from the LASA dataset were utilized, with trajectories generated from all available starting points. Prior to processing, the data underwent normalization, bringing it within the range of [$-1,1$]. The algorithm’s parameters were initialized with random values, and the maximum number of iterations was set to 2000, with a mini-batch size of 64. For the sake of equitable comparisons, DS algorithms that named SEDS and CLF-DM are chosen, for which source code was readily accessible. The parameters for these comparative algorithms were configured to match the values originally specified in their respective references, ensuring a fair benchmark.
Fig. 2 presents a visual representation of the vector fields (depicted with dark background) and the transformed trajectories generated by the proposed DS algorithm for 29 examples drawn from the LASA dataset. Notably, the reproduced trajectories (depicted in red) closely align with the original demonstrations (depicted in white). It is worth emphasizing that regardless of whether the initial point of the DS matches the starting point of the demonstrations, the reproduced DS consistently converges towards the intended goal. Furthermore, the last three images in the third row and first image in the fourth row of Fig. 2 illustrate more intricate motions. For instance, in the third image of the third row, three different types of demonstrations commence from distinct initial points but all culminate at the same target point, underscoring the algorithm’s versatility and effectiveness.
The quantified results are shown Table I, the results clearly demonstrate that the proposed algorithm outperforms other methods in terms of trajectory reproduction accuracy. Notably, the proposed algorithm achieves the highest SEA score and the best $V_{rmse}$ value, showcasing substantial improvements of approximately $15.56\%$ in SEA score and approximately $13.23\%$ in $V_{rmse}$ when compared to CLF-DM.
@@
TABLE I: Reproduction error of different DS approaches on the LASA dataset
Learning method Mean SEA($mm^{2}$) Mean $V_{rmse}(mm/s)$
@@ V CONCLUSIONS
In this paper, a neural network-based algorithm designed for learning from single or multiple demonstrations is introduced. The algorithm’s performance is assessed through both simulated scenarios featuring diverse handwriting examples and practical experiments involving a physical robot. The experimental outcomes provide compelling evidence regarding the efficacy of the proposed method.

### v6

@@ I INTRODUCTION
The popularity of robotics is growing with advancements in technology, and robots are expected to possess greater intelligence to perform complex tasks comparable to those performed by humans. However, programming them, especially for non-experts, is challenging. In such situations, Learning from Demonstration (LfD) is a user-friendly approach for robots to acquire skills by observing demonstrations, without relying on an explicit script or defined reward functions [1, 2, 3].
@@
While CLF-DM has the advantage of being able to learn from a broader set of demonstrations, it is important to highlight that the algorithm requires solving optimization problems at each step is crucial to ensuring stability. This complexity and sensitivity to parameters are important considerations.
@@
To address the challenges commonly encountered by neural network-based algorithms, a novel neural network-based approach is introduced, which eliminates the necessity of invertible transformations. This paper’s key contributions can be outlined as follows:
  • A novel neural network architecture is suggested, demonstrating the ability to directly learn a Lyapunov function from provided data.
@@
The paper is organized as follows: Section II provides a comprehensive discussion of the problem formulations. In section III, the intricate details of the neural networks are presented. In section IV, the evaluation results of the proposed algorithm. This evaluation includes simulations conducted on various handwriting trajectories sourced from the Lasa datasets [14], as well as experiments carried out on the Franka robot. The conclusion is made in section V.
@@ II Problem Formulation
where $f:\mathbb{R}^{d}\mapsto\mathbb{R}^{d}$ is a continuous and continuously differentiable function, which is described to have a single equilibrium state, denoted as $\boldsymbol{x}^{*}$. This equilibrium state represents the target state of the task and can be viewed as the attractor of the DS. To simplify the analysis and without sacrificing generality, the targets of the motions are assumed to be positioned at the origin of the Cartesian coordinate system by $\tilde{\boldsymbol{x}}=\boldsymbol{x}-\boldsymbol{x}^{*}$. Equation (1), which can also describe the manipulation skills, yields a solution $\Phi(\boldsymbol{x}_{0},t)$ that represents a valid trajectory generated when provided with $\boldsymbol{x}_{0}$. The desired motions can then be derived by incorporating this trajectory with the target state $\boldsymbol{x}=\tilde{\boldsymbol{x}}+\boldsymbol{x}^{*}$. Consequently, by altering the initial conditions $\boldsymbol{x}_{0}$, different trajectories leading to the target state $\boldsymbol{x}^{*}$ can be generated. This flexibility allows for various trajectories to be produced, each tailored to different initial conditions. The demonstration data is usually structured in the following manner: $\begin{Bmatrix}\boldsymbol{x}_{k,n},\dot{\boldsymbol{x}}_{k,n}\ |\ k=1,2,...,K_{n};\ n=1,2,...,N_{d}\end{Bmatrix}$ In this representation, $n$ denotes the index of the demonstrations, and $k$ represents the sampling time step. Here, $N_{d}$ indicates the total number of demonstrations, while $K_{n}$ signifies the total number of samples. The state variable $\boldsymbol{x}$ describes the state or configuration of either a human or a robot. The derivative $\dot{\boldsymbol{x}}$ denotes the first-order time derivative of $\boldsymbol{x}$. Additionally, it is crucial to emphasize that all demonstrations related to a specific task share the same target state. It is assumed that these demonstrations conform to an Autonomous DS as described in Equation (1). This alignment allows for the modeling of the system in a parametric fashion, which can be represented as:
@@
where $\hat{\dot{\boldsymbol{x}}}$ represents the estimated value of $\dot{\boldsymbol{x}}$, and $\hat{f}$ signifies a manually designed model aimed at approximating the DS as depicted in (1), $\boldsymbol{\theta}$ denotes the parameters within the model that are subject to learning. The optimal parameter configuration, denoted as $\boldsymbol{\theta^{*}}$, is derived through the minimization of the following objective function:
@@
where $\propto$ denotes the proportionality relation, $\left\|\cdot\right\|^{2}$ denotes the $l_{2}$-norm and $\hat{\dot{\boldsymbol{x}}}_{k,n}$ means the estimate of $\dot{\boldsymbol{x}}_{k,n}$. Hence, the objective function ensures that the learned model produces accurate reproductions closely resembling the observed behavior. Nonetheless, it’s crucial to emphasize that the stability of the model cannot be guaranteed solely by the objective function. An Autonomous Dynamical System (DS) is deemed to be locally asymptotically stable at the point $\boldsymbol{x}^{*}$ if there exists a continuous and continuously differentiable Lyapunov candidate function $V(\boldsymbol{x}):\mathbb{R}^{d}\rightarrow\mathbb{R}$ that satisfies the following conditions:
@@
In this paper, a data-driven approach utilizing neural networks to learn the Lyapunov candidate function is proposed. This method offers an alternative that bypasses the need for designing Lyapunov candidate function, facilitating the achievement of stability in the DS learning process, which is introduced in the following section.
@@ III The Proposed Approach
Figure 2: The simulation utilizing the proposed algorithm is depicted in this paper. Images with a black-yellow background display the learned vector fields. Within these visuals, the white dotted lines represent the original demonstration data, while the red solid lines depict the reproductions from identical initial points. The target points are denoted as $\cdot$” in these illustrations. Moreover, the transformed reproduction trajectories (aligned with $\boldsymbol{y}$ as illustrated in Fig. 1) are showcased against a white background. Within this context, solid points indicate the starting positions for these trajectories.
@@
$\boldsymbol{y}^{\mathrm{T}}\dot{\boldsymbol{y}}=\begin{cases}\boldsymbol{y}^{\mathrm{T}}n(\boldsymbol{y})\ \ \ \ \ \ if\ \boldsymbol{y}^{\mathrm{T}}n(\boldsymbol{y})<-\beta\boldsymbol{y}^{\mathrm{T}}\boldsymbol{y},\\ -\beta\boldsymbol{y}^{\mathrm{T}}\boldsymbol{y}\ \ \ \ \ \ \ \ \ otherwise,\end{cases}$ (13)
where $n(\cdot)$ is a neural network and $\beta$ is a manually designed small positive constant. Taking inspiration from the research conducted by [15] and simultaneously learning both $\dot{\boldsymbol{y}}$ and $\dot{\boldsymbol{x}}$, the equation (12) can be reformulated as:
$\dot{\boldsymbol{y}}=n(\boldsymbol{y})-\boldsymbol{y}\frac{\mathrm{Relu}(\boldsymbol{y}^{\mathrm{T}}n(\boldsymbol{y})+\beta\boldsymbol{y}^{\mathrm{T}}\boldsymbol{y})}{\boldsymbol{y}^{\mathrm{T}}\boldsymbol{y}},$ (14)
@@
To address the third problem, the neural network structure can be configured in a way that ensures the output becomes a zero vector only when the input is a zero vector. One simple way to enhance the performance of the neural network is to utilize a suitable activation function, such as the $\mathrm{Softplus}$ or $\mathrm{Sigmoid}$ function, which guarantee that the output is always greater than zero. Alternatively, the $\mathrm{Tanh}$ function can be employed to ensure that the output is always lesser than one. Furthermore, the output of the neural network should be multiplied with the input vector $\boldsymbol{x}$ and then added to it.
In our paper, the expression for $\boldsymbol{y}$ corresponding with the illustration in Fig. 1 that is defined as follows:
@@
Interestingly, when there is a one-to-one correspondence between the target points in two spaces, the input is $\infty$ then the output is also $\infty$, and the two spaces have the same dimensionality, the transformation can be referred to as a bijective transformation. By contradiction, the two-dimensional space is first taken into consideration, and it is assumed that the transformation from $\boldsymbol{x}$ to $\boldsymbol{y}$ is not bijective. In this case, it can be deduced that two vectors $\boldsymbol{x}$ are transformed to the same vector $\boldsymbol{y}$, such as $g(\boldsymbol{x_{1}})=g(\boldsymbol{x_{2}})$. Since equation 7 sets the Lyapunov energy function, it follows that the energy of both $x_{1}$ and $x_{2}$ is the same. In the transformed two-dimensional space, points that possess identical energy levels can be aggregated to constitute a single, closed, and continuous circle. Analogously, in the original space, points sharing the same energy level are expected to form two separate, continuous curves that intersect at points $x_{1}$ and $x_{2}$, respectively. Utilizing the Lyapunov energy function, it is established that the energy of points situated on these closed curves must be less than that of the points $x_{1}$ and $x_{2}$. Consequently, the existence of such closed curves becomes untenable, necessitating the conclusion that the two-dimensional transformation in question is indeed bijective. Furthermore, when the dimensionality of the transformation space exceeds two, a similar conclusion can be deduced by conceptualizing a closed, high-dimensional volume. Thus, it is justified to characterize the transformation as bijective across higher dimensions.
@@ IV-A Simulation Results
To assess the effectiveness of the proposed algorithm, the quantitative evaluation of accuracy was performed through the application of two error metrics: the Swept Error Area (SEA) [7] and the Velocity Root Mean Square Error ($V_{rmse}$) [16]. The SEA scores serve as an indicator of the algorithm’s capability to faithfully replicate the shapes of motions. On the other hand, the $V_{rmse}$ metric gauges the algorithm’s proficiency in preserving the velocities inherent in demonstrations. A lower SEA signifies superior accuracy in reproducing the trajectories, while a lower $V_{rmse}$ indicates that the reproductions closely emulate the smoothness observed in the original demonstrations. These two metrics comprehensively evaluate the performance of the reproductions by quantifying their resemblance to the demonstrated trajectories.
@@
For the implementation of the proposed algorithm, the ”AdamW” optimization method is opted. Specifically, the learning rate was configured at $1\times 10^{-5}$, coupled with a decay rate of $0.99$. Subsequently, the trained DS model was employed to generate trajectories starting from identical points and with matching step increments as the original demonstrations. In this paper, all data from the LASA dataset were utilized, with trajectories generated from all available starting points. Prior to processing, the data underwent normalization, bringing it within the range of [$-1,1$]. The algorithm’s parameters were initialized with random values, and the maximum number of iterations was set to 2000, with a mini-batch size of 64. For the sake of equitable comparisons, DS algorithms that named SEDS and CLF-DM are chosen, for which source code was readily accessible. The parameters for these comparative algorithms were configured to match the values originally specified in their respective references, ensuring a fair benchmark.
Fig. 2 presents a visual representation of the vector fields (depicted with dark background) and the transformed trajectories generated by the proposed DS algorithm for 29 examples drawn from the LASA dataset. Notably, the reproduced trajectories (depicted in red) closely align with the original demonstrations (depicted in white). It is worth emphasizing that irrespective of whether the starting points align with those of the demonstrations, the reproduced DS consistently converges toward the intended goal. Furthermore, the last three images in the third row and first image in the fourth row of Fig. 2 illustrate more intricate motions. For instance, in the third image of the third row, three different types of demonstrations commence from distinct initial points but all culminate at the same target point, underscoring the algorithm’s versatility and effectiveness.
The quantified results are shown Table I, the results clearly demonstrate that the proposed algorithm outperforms other methods in terms of trajectory reproduction accuracy. Showcasing substantial improvements of approximately $15.56\%$ in SEA score and approximately $13.23\%$ in $V_{rmse}$ when compared to CLF-DM.
@@
TABLE I: Variation in Reproduction Errors of different DS approaches on the LASA dataset
methods Mean SEA($mm^{2}$) Mean $V_{rmse}(mm/s)$
@@ V CONCLUSIONS
This paper introduces a neural network-based algorithm specifically designed for learning from either single or multiple demonstrations. The algorithm’s performance is assessed through both simulated scenarios featuring diverse handwriting examples and practical experiments involving a physical robot. The experimental outcomes provide compelling evidence regarding the efficacy of the proposed method.